A brass sculpture of the letter N is formed from one long metal parallelogram folded into three congruent parallelograms. What is the area of one of the smaller parallelograms? What is the total area of the long metal parallelogram?


1
Expert's answer
2020-01-20T09:39:43-0500





Given,


"base = 4 ft \\

height = 6 ft"

Area of a parallelogram

"=base \times height"


"= 4 \times 6 = 24 \space sq.ft"

From the question, all parallelograms are congruent


Area of one of the smaller parallelogram "=24 \space sq.ft"


We have 3 parallelograms.

So the total area of long metal parallelogram

"= 3 \times Area \space of \space one \space parallelogram = 3 \times 24 = 72 \space sq.ft"

Answer:

total area of long metal parallelogram ="72 sq.ft"